### 03 Feb 2024 v1

# Socialite-llama: An Instruction-Tuned Model for Social Scientific Tasks

Abstract
Social science NLP tasks, such as emotion or humor detection, are required to capture the semantics along with the implicit pragmatics from text, often with limited amounts of training data. Instruction tuning has been shown to improve the many capabilities of large language models (LLMs) such as commonsense reasoning, reading comprehension, and computer programming. However, little is known about the effectiveness of instruction tuning on the social domain where implicit pragmatic cues are often needed to be captured. We explore the use of instruction tuning for social science NLP tasks and introduce Socialite-llama— an open-source, instruction-tuned Llama2. On a suite of 20 social science tasks, Socialite-llama improves upon the performance of Llama2 as well as matches or improves upon the performance of a state-of-the-art, multi-task finetuned model on a majority of them. Further, Socialite-llama also leads to improvement on 5 out of 6 related social tasks as compared to Llama2, suggesting instruction tuning can lead to generalized social understanding. All resources including our code, model and dataset can be found through  bit.ly/socialitellama.
1 Introduction
Instruction finetuned large language models (LLMs) have demonstrated impressive performance on many standard NLP tasks Wei et al. (2022a); Chung et al. (2022), but these models tuned on non-social tasks seem to have poor social pragmatics Ziems et al. (2023); Choi et al. (2023); Havaldar et al. (2023); V Ganesan et al. (2023). The fact that instruction tuning has been successful with limited amounts of data Gupta et al. (2023) and that instruction tuned models have the ability to generalize to new tasks in both few- and zero-shot settings Wei et al. (2022a), suggests that LLMs could become more socially capable by instruction tuning them on a wide variety of social NLP tasks.
Here, we introduce Socialite-llama, a Llama2 7B-based LLM Touvron et al. (2023) instruction tuned on a suite of social scientific classification tasks spanning 5 broad categories for which we hand-craft instructions. We evaluate its performance on both zero- and few-shot settings on seen and related social tasks, demonstrating that our model significantly outperforms prior open models. Our results support prior research highlighting the effectiveness of instruction tuning when applying LLMs to a new domain.
Social and psychological factors have been shown important and beneficial to model in past interdisciplinary NLP studies Lynn et al. (2017); Flek (2020); Hovy and Yang (2021). Modeling human factors and social context can not only improve performance on primarily non-social NLP tasks  Lynn et al. (2017); Flek (2020); Hovy and Yang (2021), but can also prove to be beneficial for a number of psychological and social scientific tasks  Garten et al. (2019); Matero et al. (2021). However, a major limitation of prior models has been that they have been task-specific and do not generalize well to new tasks. Many such models also are trained from scratch and typically need an extensive amount of human-annotated training data. With the development of LLMs that exhibit the capability to learn from instructions, we posit that this new capability can be leveraged to address these two major limitations.
Our contributions include: (1) we develop and systematically evaluate Socialite-llama– an instruction-tuned language model for social science tasks – across 20 seen and 6 related social scientific tasks, (2) we show Socialite-llama consistently improved over Llama2 in all seen tasks and the improvement generalized to 5 of 6 related tasks; In fact, it matched the performance of a state-of-the-art multi-task tuned DeBERTa on a majority of seen tasks, (3) we suggest that the benefits of few-shot examples (over zero-shot) become negligible on tasks seen during instruction tuning as opposed to related tasks where few-shot still provided a benefit, and (4) we release Socialite-llama  [^1] Model available on huggingface models as well as its instructions corpora, SocialiteInstructions  [^2] Dataset available on huggingface datasets, as open-source resources for the community.
2 Related Work
Language is inherently social — the underlying meaning is constructed through social interactions Wittgenstein (1953); Clark and Schober (1992); Hovy and Spruit (2016). Understanding communication requires reasoning about the social implications drawn from that message Halliday (2004). Prior work has sought to build language models for social scientific tasks, which we discuss next.
LMs for Social Scientific NLP
Social science NLP models are usually built by fine-tuning Sap et al. (2020); Matero et al. (2019); V Ganesan et al. (2021) for specific tasks or pretraining language models on the the domain of language that captures the social factors. BERTweet Nguyen et al. (2020), the first public large-scale pre-trained language model on English tweets, was trained using the BERT architecture with the RoBERTa pre-training procedure Liu et al. (2019) for this purpose. DeLucia et al. (2022) built Bernice, a multilingual model for social science NLP tasks using the RoBERTa architecture and pretrained from scratch on 2.5 billion tweets. It outperforms a variety of models adapted to social media data as well as strong multilingual baselines. Despite the strength of such existing task-specific models, there exists no base instruction-following LLM tailored to the domain of social science tasks.
Instruction-Tuning
Instruction tuning, in general, refers to the practice of finetuning pre-trained language models to better understand and respond to a wide variety of human requests that are expressed in natural language Wei et al. (2022a); Mishra et al. (2022). The success of instruction tuning requires two key components: 1) a powerful pre-trained language model like Llama2 Touvron et al. (2023), and 2) an instruction dataset that is diverse and representative enough to adapt the LLM to potential downstream usage. However, existing instruction-tuning datasets Longpre et al. (2023); Sanh et al. (2022); Wang et al. (2023, 2022) are general-purpose and do not contain a significant amount of social science tasks. Consequently, models built using this data have limited usability for social scientific tasks Choi et al. (2023); Ziems et al. (2023). To address this gap, we curate an instruction-tuning dataset aimed at modeling social scientific knowledge and use it to train our model.
3 Datasets and Experiments
We use a diverse collection of English social science NLP tasks to create SocialiteInstructions by (re-)framing each task into a binary or multi-class classification problem. For each task, we hand-craft instructions for the model to follow. These instructions were composed with task descriptive which included examples in some cases (e.g. BiasedImplication from  Table 9), followed by the classification label choices it has to pick from. The instructions and the dataset sizes for all the tasks have been tabulated in Appendix A.
3.1 Training Tasks
We draw on datasets mentioned in Choi et al. (2023) for compiling our instruction following dataset. Our train set spans 20 different datasets across 5 broad categories of tasks: Humor, Offensiveness, Sentiment and Emotion, Trustworthiness, and other social factors. To ensure broader applicability, we frame the non-classification datasets as classification tasks. We refer to these as seen tasks going forward.
Humor We use SemEval 2021 Task 7 data Meaney et al. (2021) to capture humor. We use the binary humor detection task as is (Humor). Additionally, we convert the humor rating task into a binary classification problem collapsing ground truth labels higher than 3 into high humor, and the others into low humor (HumourRating).
Offensiveness We use Social Bias Frames Sap et al. (2020) as a benchmark to detect offense and bias directed towards people and groups. Social Bias Frames comprises of 4 binary classification tasks - Offensive to discern whether a given text exhibits rudeness, disrespect, or toxicity, Sexist to determine if a text contains lewd or sexual references which can be considered as offensive, IntentToOffend to capture whether the perceived motivation of the author was to indeed offend, and BiasedImplication to identify any forms of prejudice or group-based discrimination within the text.
Sentiment and Emotion For emotion classification, we use SemEval 2018 Task 1 data Mohammad et al. (2018). Emotion involves classifying a tweet into anger, joy, optimism or sadness displayed by the author. We also include Emobank Buechel and Hahn (2017) to infer the valence, arousal and dominance levels that a text would invoke in a reader. We convert these regression tasks to binary classification - by transforming a score greater than 4 to ’high’ and less than 4 to ’low’, for each of the three datasets respectively (ValenceCLS, ArousalCLS, DominanceCLS). We use SemEval 2017 Task 4 data for sentiment classification Rosenthal et al. (2017) to identify the overall sentiment of a text as positive, negative or neutral. Further, we also include SameSideStance Körner et al. (2021), the task of classifying whether two different texts are on the same side of an argument.
Trustworthiness We use Hyperbole Zhang and Wan (2022) to detect hyperbolic language present in a text, and SubjectiveBias Pryzant et al. (2020), to detect which sentence exhibits subjective bias among two pieces of text.
Other Social Factors We also include other discursive and rhetorical type tasks — empathy scored on multi-item scales (EmpathyExplorations) Sharma et al. (2020), self-rated empathy (EmpathySelfRated) Buechel et al. (2018), distress (DistressSelfRatedBuechel et al. (2018), figurative speech detection (FLUTEChakrabarty et al. (2022), politeness (PolitenessHayatiHayati et al. (2021) and intimacy (IntimacyPei and Jurgens (2020).
3.2 Evaluation Tasks
For evaluation, we use the same twenty seen datasets as well as six datasets for related social tasks. The process we used to choose the ‘seen tasks’ for instruction tuning was towards the goal of training on a very broad category of tasks and consequently, any task would be related to these fundamental categories seen during training. Hence for evaluation we choose six tasks, half of which are more directly related to seen tasks than the other. The datasets used for evaluation contained no data overlap with the training tasks’ data. Among the related tasks’ datasets, we first select three datasets directly related to our training tasks — HateSpeech Basile et al. (2019) to detect hate speech in a tweet, PolitenessStanford Fu et al. (2020) to judge whether a text is polite or impolite, and AgreeDisagree Varadarajan et al. (2022) to classify whether two texts agree, disagree or N/A on a particular topic. To strengthen our evaluation, we also evaluate on three other tasks which are less related to the seen set — Irony Van Hee et al. (2018) to classify if a tweet is ironic, Optimism Ruan et al. (2016) to categorize a text as optimistic, pessimistic or neutral, and Complaints Preoţiuc-Pietro et al. (2019) to judge if a text is a complaint or not. We refer to these six as related social tasks.
3.3 Task Selection Criterion
The above tasks were chosen from SocKET (a) to be representative of each social category and (b) if the document lengths weren’t too long, i.e., utterance level tasks Ziems et al. (2023).
Our initial experiments suggested that skew in sample sizes from different tasks affected the overall performance. Particularly, using all 35k examples from Social Bias Frames for all the four tasks limited the performance on PolitenessHayati and SameSideStance. Hence, we picked datasets from SocKET based on 2 additional factors: recency and minimum number of 8k examples. This is why we chose Sap et al. (2020) over other toxicity datasets (further down sampled 8k examples for each task), and Rosenthal et al. (2017) over Socher et al. (2013) for sentiment task.
Task
Socialite-
Llama
Llama2
Few-shot Zero-shot Few-shot Zero-shot
Seen Tasks
Sentiment 72.69 73.43 46.40 18.28
Emotion 79.68 81.08 66.84 23.92
ValenceCLS 82.34 83.84 44.52 31.59
ArousalCLS 60.69 64.22 52.69 34.42
DominanceCLS 60.88 66.63 43.05 19.33
EmpathyExplorations 30.10 56.99 29.43 5.89
EmpathySelfRated 63.59 63.97 58.99 8.75
DistressSelfRated 68.45 70.96 42.36 9.80
FLUTE 80.77 96.27 55.13 4.81
Hyperbole 65.62 69.23 59.20 25.43
SameSideStance 88.19 88.19 82.10 25.00
Humor 90.48 95.29 44.44 38.8
HumourRating 52.81 48.66 43.51 34.84
PolitenessHayati 89.36 84.54 83.58 24.34
Intimacy 26.25 33.21 13.24 3.17
SubjectiveBias 52.04 87.13 51.18 40.98
Offensive 80.47 83.11 71.78 35.76
Sexist 72.61 73.85 52.81 6.25
IntentToOffend 74.17 75.65 32.09 21.34
BiasedImplication 79.66 85.72 56.41 26.16
Related Tasks
HateSpeech 42.13 39.63 68.66 31.34
Irony 42.99 59.04 47.11 32.24
PolitenessStanford 63.22 64.89 58.64 1.2
Optimism 62.24 36.96 56.61 44.49
Complaints 79.35 85.37 58.23 22.8
AgreeDisagree 39.54 19.3 28.1 10.01
Table 1: Evaluation of Socialite-llama and Llama2 on 20 seen tasks and 6 unseen tasks using macro-F1 scores. We report few-shot performance with $k=5$ for Socialite-llama on seen tasks and $k=15$ on unseen tasks. For Llama2, we use $k=15$ for all tasks since it is not trained for social scientific tasks. $k$ refers to the number of examples used in the few-shot setting. All differences between Socialite-llama and Llama2 are statistically significant $(p<0.001)$. Bold indicates the performance with the best setting for that task.
3.4 Training Procedure
We transform each of the 20 training datasets using hand-crafted instructions. To reduce the skew caused by the disproportionate number of examples present in a few tasks, we limit the number of training examples from Offensive, Sexist, IntentToOffend, BiasedImplication, Sentiment and SubjectiveBias to 8k and use the original number of training data points for the rest. This resulted in ~108k data points. We employ Low-Rank Adaptation Hu et al. (2022), with rank set to 8 applied on the key and query projection matrices reducing the number of trainable parameters to 4.1 million, and use AdamW Loshchilov and Hutter (2019) with a learning rate of $\eta=1e-4$ and a batch size of $64$ for a maximum of $7$ epochs with early stopping. We expand the context length to 3k using mixed precision training, gradient checkpointing and DeepSpeed Zero Rajbhandari et al. (2020) Stage 2 which enables finetuning on 2 48GB A6000 GPUs in ~21 hours and train using Huggingface Wolf et al. (2020). Training Socialite-llama on Lambda [^3] https://lambdalabs.com would cost \$33.60, making it very cost-effective.
4 Results
We analyze the effect of instruction tuning Llama2 on 20 seen and 6 related social scientific tasks and report macro F1 scores in Table 1.
Instruction tuning improves social understanding. We show that Socialite-llama outperforms Llama2 on all seen tasks in both zero- and few-shot setting. In fact, Socialite-llama used in a zero-shot setting is even better than Llama2 in few-shot setting. Surprisingly, we find that the benefit of few-shot learning over zero-shot is much less on seen tasks as opposed to the related tasks with the instruction tuned model.
Particularly, performance of our model is better for all tasks that are proximate to personal factors or how one feels, such as the affective dimensions, sentiment, emotion and empathy/distress. We also see consistent improvements over Llama2 on offensiveness related tasks. Despite the tremendous general-purpose capabilities of Llama2, it appears to lack social understanding from text, and can be improved using instruction tuning on social scientific tasks.
Task
Socialite-
LLama
DeBERTa
(multi-task FT)
Sentiment 73.43 69.00
Emotion 81.08 80.00
DistressSelfRated 70.96 65.00
EmpathySelfRated 63.97 59.00
Hyperbole 69.23 69.00
SameSideStance 88.19 76.00
Humor 95.29 91.00
PolitenessHayati 84.54 89.00
Intimacy 33.21 46.00
SubjectiveBias 87.13 96.00
Offensive 83.11 83.00
IntentToOffend 75.65 74.00
Sexist 73.85 79.00
BiasedImplication 85.72 87.00
[…]
Instruction tuning helps generalization. In related social science tasks, which are composed of new instructions and language samples, we find that Socialite-llama is better on 4 out of 6 under few-shot setting and 5 out of 6 overall. These 6 related tasks were on a varying degree of relatedness to the ones in the seen tasks. Socialite-llama shows significant improvement in performance over Llama2 on the related social science tasks. Through instruction tuning, our model generalizes social aspects of language to perform well on new related social tasks and task categories.
To assess the generalization emerging from the commonalities present between the task, we applied the instructions from the seen tasks’ set on a related task and compared its performance with the related tasks’ instruction (which was unseen to Socialite-llama). A non-author NLP expert suggested 4 seen tasks (Offensive, Sexist, IntentToOffend and BiasedImplication) that they thought were very similar to the HateSpeech task. Of those, instructions from Offensive and IntentToOffend gave improvement over zero-shot HateSpeech performance, while the other two led to worse performance. This suggests it’s possible to do better on zero-shot by using a related task prompt from instruction tuning but it can also hurt performance.
Socialite-llama is state-of-the-art overall on SocKET. We compare Socialite-llama with a state-of-the-art multi-task finetuned DeBERTa model presented in Choi et al. (2023) for a subset of the seen tasks since their DeBERTa model can’t be applied on the related social tasks. Despite being trained on significantly less data than DeBERTa model (2.1 million data points and 58 tasks as compared to 108k data points and 20 tasks), Table 2 shows that our model is better or equivalent in performance on 10 out 14 tasks. This result highlights a large reduction in the performance separation between LLMs and smaller task finetuned language models in Social Science tasks established by previous works Ziems et al. (2023); Choi et al. (2023).
5 Conclusion
We introduce Socialite-llama, an instruction tuned Llama2 7B for social science applications. Socialite-llama is trained using SocialiteInstructions, a diverse collection of 20 social scientific tasks and tested on those plus an additional 6 related social tasks. It consistently performs better than Llama2 on all the seen tasks and all but 1 related task, demonstrating its strong generalization abilities. Instruction tuning on diverse social scientific tasks helps our model achieve state-of-the-art overall. Our results provide valuable insights to improve the social understanding of LLMs.
Limitations
The improvements brought through instruction tuning of Llama2 on social scientific tasks highlights the prevailing neglect of these tasks and the importance of doing this. However, there are some important limitations. For one, the tests for generalization to unseen tasks were limited to those that are semantically related to seen tasks (many tasks in SocKET are related to each other). Future work would need to evaluate this model and process for generalizing to more distant tasks as well as across other points in time.
This work was informed by a number of works in the NLP and deep learning literature, was largely facilitated by various open sourced datasets Choi et al. (2023) and code Wolf et al. (2020). However, the number of tasks in our work is not close to the number of tasks found in the prior instruction tuning literature Wang et al. (2022); Sanh et al. (2022). This work would likely benefit more tasks within the realm of social science NLP.
Owing to limited access to computational resources, we could only focus on classification tasks on the smallest available Llama2 model and Low-Rank Adaptation as training strategy. However, literature on scaling models Brown et al. (2020); Wei et al. (2022b) suggests that these performances likely go up with increasing the model size, datasets, instructions, which would make it more capable for regression and generation tasks with increased input lengths. Inclusion of few-shot examples in the training set could also improve the few-shot generalization capability of the model, which could not be performed with the resources available.
Ethics Statement
Large language models (LLMs) like GPT-4 OpenAI (2023) have been deployed to tens of millions of consumers in different forms Heaven (2023); Hu (2023). While a number of works have already shown its limited social understanding Ziems et al. (2023); Choi et al. (2023); Havaldar et al. (2023); V Ganesan et al. (2023), their rapid adoption and proliferation necessitates improving their social sensibility to make it reliable and useful. This work is aimed towards building a socially sensible language model for the practitioners of social science NLP research.
We hope our work improves NLP’s benefit for social scientific pursuits and caution against the use of such models for non-scientific pursuits like targeting of individuals without their consent or awareness. This work has been built for research and can be re-distributed as per the policy laid out by Meta for Llama2  [^4] https://ai.meta.com/llama/license/. We strongly believe that these models have to be tested for failure modes and harmful biases, and should be further be adjusted through appropriate methods Ouyang et al. (2022) before deploying it into the hands of users.
Acknowledgements
We would like to thank the reviewers for their valuable feedback that helped us improving this paper. This work was supported in part by a grant from the CDC/NIOSH (U01 OH012476), a grant from IARPA (585968), a grant from the NIH-NIAAA (R01 AA028032). YKL was supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the NSF under the award IIS #2007290. We approve this work for public release and redistribution with attribution. The conclusions and views in this paper are those of the authors, and should not be seen as representing the official policies expressed or implied by CDC, NIOSH, NIH, NIAAA, IARPA, DARPA, ARFL, NSF or the U.S. government.
Appendix A Appendix
We first provide statistics about SocialiteInstructions which has been used to instruction tune Socialite-llama. Table 3 includes information about amount of data used per task as well as the number of labels used.
Next, we provide descriptions of the instructions and prompt templates used for each task in SocialiteInstructions. Please note that this may contain examples of potentially dangerous and harmful text.
Dataset Train Set Validation Set Test Set Num_classes
Emotion 3257 374 1421 4
FLUTE 6780 754 1498 4
EmpathyExplorations 2220 247 617 3
Humor 8000 1000 1000 2
Offensive 8000 4666 4691 2
Sexist 7999 4666 4691 2
IntentToOffend 7999 4666 4691 2
BiasedImplication 7999 4666 4691 2
PolitenessHayati 256 32 32 2
Hyperbole 2580 323 323 2
SameSideStance 140 18 17 2
Sentiment 8000 2000 12284 3
Intimacy 1797 225 225 6
SubjectiveBias 8000 9379 9379 2
ValenceCLS 9002 510 550 2
ArousalCLS 9002 510 550 2
DominanceCLS 9002 510 550 2
EmpathySelfRated 1487 186 186 2
DistressSelfRated 1487 186 186 2
HumourRating 4932 632 615 2
HateSpeech - - 2970 2
Irony - - 784 2
PolitenessStanford - - 567 2
Optimism - - 1495 3
Complaints - - 345 2
AgreeDisagree - - 4760 3
Table 3: Training, validation and test set statistics SocialiteInstructions. ‘-’ denotes that the dataset was not used in training Socialite-llama but we create instructions for its test set. Overall, SocialiteInstructions contains ~202k data points.
Dataset Instruction Format Example
Humor
Instruction: Upon receiving a piece of text, your task is to analyze and determine whether the intention of the text was to be humorous. You are instructed to look at the text and identify the structure of the joke, e.g. setup and punchline, or the content of the joke, e.g. absurdity, in order to determine if the intention of the text was to be humorous. If you think the intention of the text was to be humorous, classify it as ‘humorous’, else classfy it as ‘non-humorous’.
Input: TENNESSEE: We’re the best state. Nobody even comes close. *Elevennessee walks into the room* TENNESSEE: Oh shit…
Output: humorous
HumourRating
Instruction: Upon receiving a piece of text, your task is to assess its comedic quality and categorize it as either ‘low humor’ or ‘high humor’.
Input: How many men does it take to open a can of beer? None. It should be open by the time she brings it to the couch.
Output: low humor
Table 4: Instruction prompts and output examples for humour task
Dataset Instruction Format Example
Sentiment
Instruction: Evaluate the sentiment conveyed in the input text and determine whether it is positive, negative, or neutral. This sentiment assessment should encompass the overall sentiment of the event described within the context of the topic mentioned in the text. Your options for classification are confined to positive, negative or neutral.
Input: Few more hours to iPhone 6s launch and im still using the 4th generation
Output: positive
Emotion
Instruction: Analyze the following sentence and determine the predominant emotion it displays. Your options for classification are confined to anger, joy, optimism, or sadness. Please select one emotion from the given alternatives that you believe best epitomizes the emotional context of the sentence.
Input: Worry is a down payment on a problem you may never have’. Joyce Meyer. #motivation #leadership #worry
Output: Optimism
ValenceCLS
Instruction: Analyze the provided text using the Valence-Arousal-Dominance model for emotional assessment. Your task is to classify the valence level it would likely elicit in an average reader, where ‘Low Valence’ indicates a low level of pleasant feelings and ‘High Valence’ indicates a high level of pleasant feelings. Remember, the valence scale is used to measure the degree of pleasure or displeasure a person may feel towards something. Your options for classification are confined to ‘Low Valence’ or ‘High Valence’.
Input: Remember what she said in my last letter?
Output: High Valence
ArousalCLS
Instruction: Analyze the provided text using the Valence-Arousal-Dominance (VAD) emotional model. Your task is to classify the arousal level it might trigger in an average reader. Arousal, in this context, refers to the degree of energy or lethargy the text might induce. ’Low Arousal’ indicates a low arousal level, suggesting the text is likely to make the reader feel calm or lethargic. Conversely, ’High Arousal’ indicates a high arousal level, suggesting the tweet is likely to energize or excite the reader. Your options for classification are confined to ‘Low Arousal’ or ‘High Arousal’.
Input: Remember what she said in my last letter?
Output: High Arousal
DominanceCLS
[…]
Input: Remember what she said in my last letter?
Output: High Dominance
SameSideStance
Instruction: You are provided with two pieces of text sourced from an online debate forum. Your task is to analyze and categorize these texts based on their argumentative stance. Determine whether both texts are arguing in favor of the same viewpoint or if they are opposing each other. Your options for classification are confined to ‘same side’ or ‘not same side’.
Input: Legalizing gay marriage will not destroy man/woman relationships. [SEP] Why should gay marriage destroy man/woman relationships? It would just give gays the option to marry.
Output: same side
Table 5: Instruction prompts and output examples for sentiment and emotion task
Dataset Instruction Format Example
SubjectiveBias
Instruction: Given two pieces of text, your objective is to detect subjective bias, which manifests when language that should remain neutral and impartial is influenced by feelings, opinions, or personal preferences, whether intentionally or unintentionally. If you find bias in the first sentence, indicate ‘first sentence’ as the output; otherwise, specify ‘second sentence’.
Input: the term finds widespread usage among members of the educational establishment who see students as tools of social change. [SEP] the term finds widespread usage among members of the educational establishment who see students as agents of social change.
Output: first sentence
Hyperbole
Instruction: Upon receiving a piece of text, your task is to analyze and determine whether it contains hyperbolic language, which is an exaggerated statement or claim not meant to be taken literally, or if it does not. Your options for classification are confined to ‘hyperbole’ or ‘not hyperbole’.
Input: He looks great and yet he must be pushing sixty by now.
Output: not hyperbole
Table 6: Instruction prompts and output examples for trustworthiness task
Dataset Instruction Format Example
EmpathyExplorations
Instruction: Evaluate the degree of inquiry exhibited in the counselor’s response provided below, categorizing it as either “Strong Exploration", “Weak Exploration" or “No Exploration". We define ‘exploration’ as instances where a mental health counselor displays keen interest in a patient by asking about experiences that haven’t been explicitly mentioned.
Input:
Patient: I’m finally on medication and found a therapist who makes me feel hopeful for the first time in years. It has been a while since I’ve felt sincere hope. Counselor’s response: im envious of the fact that you want hope.
Output: No Exploration
EmpathySelfRated
Instruction: Please carefully peruse the subsequent text, which is a personal account penned by an individual expressing their emotions and reflections after reading a news article. This account is directed towards their friends. After reading, your task is to accurately classify the level of empathetic concern demonstrated by the author. Your options for classification are ‘low empathy’ which indicates low empathetic concern or ‘high empathy’ which signifies a high degree of empathetic concern.
Input: This sounds like a horrible accident. I can’t even imagine what that family is going through. The ones that were involved in the accident but survived probably saw some horrible things, and those kids who witnessed their mother die. As a parent myself that’s a terrifying thought I would never want my son to go through something like that.
Output: high empathy
DistressSelfRated
Instruction: Please carefully peruse the subsequent text, which is a personal account written by an individual to their friends. This account details their emotional reactions and cognitive responses upon reading a specific news article. Your task is to accurately classify the level of personal distress experienced by the author. Your options for classification are ‘low distress’ or ‘high distress’.
Input: This sounds like a horrible accident. I can’t even imagine what that family is going through. The ones that were involved in the accident but survived probably saw some horrible things, and those kids who witnessed their mother die. As a parent myself that’s a terrifying thought I would never want my son to go through something like that.
Output: low distress
FLUTE
Instruction: Please follow these steps:
1. First, you’ll be presented with a premise and a hypothesis in the input section.
2. Your task is to determine and categorize the type of figurative language utilized in the hypothesis.
3. Finally, based on your assessment, respond with a single answer that most accurately represents the figurative language detected in the hypothesis. Choose from these four classifications: Idiom, Metaphor, Sarcasm, or Simile.
Input:
Premise: he was flying high happy and joy.
Hypothesis: he was feeling down hurt and sad.
Output: Metaphor
Table 7: Instruction prompts and output examples for other social factors
Dataset Instruction Format Example
PolitenessHayati
Instruction: Upon receiving a piece of text, your task is to analyze and determine whether the language used within it is courteous and respectful, indicating politeness, or if it contains disrespectful or rude elements, indicating impoliteness. Your options for classification are confined to ‘impolite’ or ‘polite’.
Input: @selfiequeenbri: cause I’m tired of you big bitches coming for us skinny girls!!
Output: impolite
Intimacy
Instruction: Assess the degree of intimacy expressed in the input text, taking into account the social context within the text. Your options for classification are confined to ‘very intimate‘, ‘intimate’, ‘somewhat intimate’, ‘not very intimate’, ‘not intimate’ or ‘not intimate at all’.
Input: What are the most mediocre animals in the animal kingdom?
Output: not intimate
Table 8: Instruction prompts and output examples for other social factors
Dataset Instruction Format Example
Offensive
Instruction: Evaluate the given text for any offensive content, which includes rudeness, disrespect, or toxicity. This assessment should consider if the text could potentially offend anyone, based on previous studies indicating a higher recall rate. Identify any disrespectful, inappropriate, or harmful language, phrases, or sentiments. If these elements exist, label the text as ‘offensive’. If these elements are absent, mark the text as ‘not offensive’.
Input: jack and jill went up a hill . <repeat> jill came down , i went up , and blew jack . ooohh ! <allcaps> it do n’t rhyme . it just happened that way.
Output: not offensive
Sexist
Instruction: Analyze the provided sentence and evaluate if it contains any elements that could be considered as gender-based discrimination. You are required to categorize the sentence into one of two classifications: ‘sexism’ if it exhibits gender discrimination, or ‘not sexism’ if it does not.
Input: "Alternative Theory: They had to change the timeline to explain why the building engineer was up there–had he heard shots, he would not have gone to 32nd floor– in order to have #StephenSchuck corroborate #JesusCampos story, it had to have been before shooting started.
Output: not sexism
IntentToOffend
Instruction: Your task is to analyze the given text and identify if it intentionally promotes social biases and stereotypes. If you perceive that the author’s motive was to deliberately offend, categorize the text as ‘intentional’. If there’s no apparent intent to offend, classify it as ‘not intentional’. Remember, the focus is on the author’s intent to offend, not just the offensiveness of the content.
Input: should i start discourse with eileen and hope it makes them stop telling my friends im a pedophile cus i’m fucking not?
Output: not intentional
BiasedImplication
Instruction: Your task is to scrutinize a given text and identify any signs of prejudice. You need to differentiate between statements that merely insult an individual and those that imply group discrimination, such as comparing ‘F*ck you’ (an individual insult) with ‘F*ck you, f*ggot’ (a group implication). Based on your analysis, categorize the text as either ‘biased’ or ‘not biased’.
Input: Obviously everyone you disagree with on twitter is a manipulation bot & everyone you agree with is real, right? :p
Output: biased
Table 9: Instruction prompts and output examples for offensive task
Dataset Instruction Format Example
HateSpeech
Instruction: Analyze the given text for hateful contents against women or immigrants. Your options for classification are confined to ‘hate speech’ or ‘not hate speech’.
Input: Send home migrants not in need of protection, Peter Dutton tells UN, HEY DUTTON HOW ABOUT THE ONES THAT HAVE STAYED AND NOT LEFT THE COUNTRY WHEN THEY SHOULD OVERSTAYERS ? WHY DONT YOU GO AND ROUND ALL THEM UP ?
Output: hate speech
Irony
Instruction: Classify the input sentence as ironic or not ironic. Here ironic can be defined as the use of words or situations to convey a meaning that is opposite to what is expected. Your options for classification are confined to ‘ironic’ or ‘not ironic’.
Input: Now I remember why I buy books online @user #servicewithasmile
Output: ironic
PolitenessStanford
Instruction: Analyze the provided text, considering its tone and language, and categorize it as either polite or impolite. Your options for classification are confined to ’impolite’ or ’polite’.
Input: Hi, I made on comment on the image page about this being different than Sbarro3 image. Please… comment?
Output: polite
Optimism
Instruction: Analyze the sentiment of the provided text and classify it as ‘optimistic’, ‘pessimistic’ or ‘neutral’.
Input: fuck you fuck you fuck you
Output: pessimistic
Complaints
Instruction: Given an input text, identify if it contains a complaint or not. Complaining is a basic speech act used to express a negative mismatch between reality and expectations towards a state of affairs, product, organization or event. Your options for classification are confined to ‘complaint’ or ‘not complaint’.
Input: @SamsungSupport Can someone please help me? I’ve already sent a DM.
Output: complaint
AgreeDisagree
Instruction: You are provided with two pieces of text and your task is to analyze and categorize these texts based on their argumentative stance. Determine whether both texts are arguing in favor of the same viewpoint, if they are opposing each other or if they are talking about two different topics. Your options for classification are confined to ‘agree’, ‘disagree’ or ‘N/A’.
Input: President Trump also intentionally lied to hide the truth from the American people and make himself look better. [SEP] President Trump told public lies every single day for his first 40 days.
Output: agree
Table 10: Instruction prompts and output examples for related social tasks tasks